OpenStudy (anonymous):

Is it the same formula? @jdoe0001

OpenStudy (anonymous):

convert the interest to annually then use pert

OpenStudy (jdoe0001):

ahmm nope... this is a compounded interest acccoiunt, and is compounded monthly meaning there are 12 months in 1 year.. so n = 12 \(\bf A=P\left(1+\frac{r}{n}\right)^{nt} \\ \quad \\ A=\textit{current amount}\\ P=\textit{original amount deposited}\to \$742\\ r=rate\to 3.4\%\to \frac{3.4}{100}\to 0.034\\ n=\textit{times it compounds per year}\to 12\\ t=years\to 15 \\ \quad \\ A=742\left(1+\frac{0.034}{12}\right)^{12\cdot 15}\)

OpenStudy (anonymous):

1234.755

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

That would be the final answer? (In other words, no more steps?)

OpenStudy (jdoe0001):

yeap..that'd be it, the amount accrued after 15 years of compounding monthly at 3.4%

OpenStudy (anonymous):

Thank you! :D